OpenStudy (anonymous):

An 400 ohm resistor contains a silver band which means it has a 10% tolerance factor. Find the upper and lower limits of the 400 ohm resister.

OpenStudy (anonymous):

10%of 400=10/100 *400=40 thus upper limit is 400+40=440 and lower limit is 400-40=360

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome dear

OpenStudy (anonymous):

i've got another if you dont mind taking a look at

OpenStudy (anonymous):

if i can i will help..

OpenStudy (anonymous):

Eric can clean the car in 24 minutes. How long does it take Jackie to wash the car alone if Eric and Jackie can clean the car together in 8 minutes?

OpenStudy (anonymous):

here in this type o f sums u hav to find that in unit time what part of the work is done

OpenStudy (anonymous):

in 1 min Eric completes 1/24 of work also in 1minute both of them together complete 1/8 of the work

OpenStudy (anonymous):

thus in 1 min Jackie completes 1/8-1/24 of the work now 1/8-1/24=(3-1)/24=2/14=1/12 thus Jackie alone can complete the work in 12 mins

OpenStudy (anonymous):

hope its clear to you

OpenStudy (anonymous):

nice! i think i got it

OpenStudy (anonymous):

welcome